OpenStudy (anonymous):

7^x = 124 Help! I have an entire homework sheet of these things and it makes no sense to me =(

OpenStudy (anonymous):

me 2 i hate homework :\

OpenStudy (anonymous):

Change of base formula log124/(log7) = 2.48

OpenStudy (anonymous):

a^x = b x = log(b)/loga every time.

OpenStudy (anonymous):

\[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (rajee_sam):

satellite always goes Au Naturale.. only natural logs whenever he answers

OpenStudy (anonymous):

ok so what happens when they give you one of these things? 3log6 (x+1) = 9

OpenStudy (rajee_sam):

asking who?

OpenStudy (anonymous):

anyone.

OpenStudy (rajee_sam):

is it base 6?

OpenStudy (anonymous):

yes

OpenStudy (rajee_sam):

\[\log_{6} (x + 1) = 9/3 = 3\]

OpenStudy (rajee_sam):

(x+ 1) = 6³

OpenStudy (rajee_sam):

x+ 1 = 216; x = 215

OpenStudy (anonymous):

woah! thank you! That makes sense. So the next one log(base)5 x = 3 does that become x = 3^5 x = 243 ?

OpenStudy (anonymous):

ah! 5^3 nvm

OpenStudy (anonymous):

Ok I'm lost again. log(base5) 2x-5 = -4

OpenStudy (rajee_sam):

5^ -4 = 2x - 5

OpenStudy (rajee_sam):

what?

OpenStudy (rajee_sam):

that's not right

OpenStudy (whpalmer4):

Whoops, mistyped...

OpenStudy (whpalmer4):

raising the base to both sides: \[5^{\log_5(2x-5)} = 5^{-4}\]\[2x-5=\frac{1}{5^4}\]\[2x=5+\frac{1}{5^4}\]\[x=\frac{1}{2}(5+\frac{1}{5^4}) = \frac{1563}{625}\]

OpenStudy (rajee_sam):

Instead of touching your nose straight from the front you are touching it from around your head.

OpenStudy (whpalmer4):

Mathematician reducing everything to a previously solved problem.

OpenStudy (rajee_sam):

\[\log_{a} B = X \space means \space a ^{X} = B\]

OpenStudy (whpalmer4):

Different people find different explanations of equivalent statements useful.

OpenStudy (anonymous):

woah gimme a sec here. I'm getting lost in a sea of numbers and letters.

OpenStudy (anonymous):

ok I think I got it (thank you guys btw!) So 3 (5)^2x-3 = 6 5^2x -3 = 2 2x -3 = log2 /log 5 am I on the right track yall?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then add three and divide by 2 and you are done

OpenStudy (anonymous):

PRAISE JESUS!!!! I love you people. Seriously like half of my GPA is due to you guys.